OpenStudy (anonymous):

Write the answer in both standard and slope intercept form. contains (-2,3) (1,4) ?

OpenStudy (anonymous):

\[y-y_1 = m(x-x_1)\]is standard form

OpenStudy (anonymous):

\[y=mx+b\]is slope intercept form, where m is slope and b is the y-intercept

OpenStudy (anonymous):

can you explain that? what numbers would i use to plug in?

OpenStudy (anonymous):

\[m=\frac{ y_2-y_1 }{ x_2-x_1 }\] use this equation to find the slope first you have 2 sets of points (x1, y1) and (x2, y2) which are (-2,3) (1,4) in this case

OpenStudy (anonymous):

find your slope, then use the standard form to find an equation. then convert it to slope intercept form

OpenStudy (anonymous):

so: 4-3 over 1- -2 = m

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1/3 should be your slope

OpenStudy (anonymous):

yep that';s what i got. so what would i use to plug in for my slope intercept: y=mx+b? i know m is 1/3

OpenStudy (anonymous):

well lets do standard form first find the standard form equation, and then use that to solve for y by itself, which should give you the slope intercept form

OpenStudy (anonymous):

Ax+By=C

OpenStudy (anonymous):

that is the general form, not standard form

OpenStudy (anonymous):

y-y1=m(x-x1) ?

OpenStudy (anonymous):

yes, use the m that you found and your points plug them in and you have standard form of the line

OpenStudy (anonymous):

y-3=1/3(x- -2)

OpenStudy (anonymous):

yes, that is the equation for the line in standard form

OpenStudy (anonymous):

now solve for y

OpenStudy (anonymous):

you will have to distribute the 1/3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

y=1/3x^2 + 2/3x +3

OpenStudy (anonymous):

no, where did you get x^2 from?

OpenStudy (anonymous):

remember, you're multiplying by just 1/3, not 1/3x

OpenStudy (anonymous):

1/3x(x+2) i distributed 1/3x to the x. (1/3*x) and then did 1/3*2

OpenStudy (anonymous):

it is not 1/3x(x+2) it is 1/3(x+2) you're adding an x that is not there

OpenStudy (anonymous):

oh. i get it. so y=1/3x + 2/3x + 3 and then i would add 1/3x to 2/3x and get 1 so y=x+3 ?

OpenStudy (anonymous):

you're distributing wrong 1/3(x+3) becomes 1/3x +2/3 y-3 = 1/3x+2/3 y = 1/3x+2/3+3 y = 1/3x +11/3

OpenStudy (anonymous):

oh. so that is the standard form?

OpenStudy (texaschic101):

m = (y2 - y1) / (x2 - x1) ...set 1 (-2,3) set 2 (1,4) m = (4 - 3)/ (1 - (-2) m = 1/3 slope intercept form : y = mx + b using slope 1/3 and either point....I will use (1,4) y = mx + b 4 = 1/3(1) + b 4 = 1/3 + b 4 - 1/3 = b 12/3 - 1/3 = b 11/3 = b so your slope intercept form of the equation becomes : y = 1/3x + 11/3 Standard form : Ax + By = C y = 1/3x + 11/3 -1/3x + y = 11/3 (multiply by 3) -x - 3y = 11 (now multiply by -1 to make x positive) x + 3y = -11 so your standard form is : x + 3y = - 11

OpenStudy (texaschic101):

do you have any questions ?

OpenStudy (texaschic101):

oops...I messed up standard form y = 1/3x + 11/3 -1/3x + y = 11/3 (multiply by 3) -x + 3y = 11 (now multiply by -1 to make x positive) x - 3y = - 11 Thats better :)